OpenStudy (anonymous):

Use the arc length formula and the given information to find s. Show your work for full credit. r = 20 ft θ = 19° s = ?

OpenStudy (anonymous):

What is s?

OpenStudy (anonymous):

thats what i have to find

OpenStudy (anonymous):

19 deg=19pi rad s=20*19pi

OpenStudy (anonymous):

thats where i got to.. im not sure where to go from this point

OpenStudy (anonymous):

Oh, its the arc lenght ok. The radian transformation is not 19pi. Its \[\frac{19 \times \pi}{180}\]Then the arc lenght is 20 times that.

OpenStudy (anonymous):

wait.. im comfused

OpenStudy (anonymous):

When you transform degrees in radians (because the formula r angle just works with radians) you use pi/180 because pi is the equivalent of 180 with radians, if you divide pi by its equivalent in degrees, you will divide the angle in 180 angles, and each one of those is one degree, but showed in radians because we divided pi, now if you have 19 degrees, and you know that one degree in radians is pi/180 you just multiply to get 19 in radians too.

OpenStudy (anonymous):

so its not 20(19pi) itd be 19 x pi / 180 ??? but what would i do from there?

OpenStudy (anonymous):

No, its 20(19pi/180)

OpenStudy (anonymous):

That is s already isnt that what you want?

OpenStudy (anonymous):

yah but i dont have to solve it?

OpenStudy (anonymous):

thatd be the final answeR?

OpenStudy (anonymous):

Yes, you do 19*(pi/180) to get the angle in readians and then you multiply that by 20 to get the lengh.

OpenStudy (anonymous):

so for the answer would i just put .. 20(19pi/180) ??

OpenStudy (anonymous):

0.105*pi multiplied 20 = 2.10pi

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Yes, thats it

OpenStudy (anonymous):

ok thank you!!